# Liquidity fragmentation on decentralized exchanges

### v5

@@ 2.1 Equilibrium
First, consider the liquidity traders’ decisions. Faced with pool sizes of $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ in the low and high pool respectively, they choose the pool which minimizes their trading costs. Conditional on trading, the small liquidity takers choose the $\ell$ fee pool. Thus, small traders arrive at the $\ell$ fee pool at rate $\theta$, and each trade one unit. By assumption the large liquidity taker wants to trade more than the posted liquidity and exhausts both pools. Thus, liquidity is consumed on the low fee pool at rate $\frac{\theta}{\mathcal{L}_{\ell}}\mathop{}\!\mathrm{d}t$, while liquidity is only consumed on the high fee pool if a large trader arrives. Once each pool is empty, liquidity providers refill it and restart the cycle. Let $d_{k},k=\ell,h$ denote the duration of a liquidity cycle on the low and high pool respectively. Then, the expected duration of a cycle on the low fee pool is
$d_{\ell}$ $=e^{-\lambda\frac{\mathcal{L}_{\ell}}{\theta}}\frac{\mathcal{L}_{\ell}}{\theta}+\int_{0}^{\frac{\mathcal{L}_{\ell}}{\theta}}t\lambda e^{-\lambda t}\mathop{}\!\mathrm{d}t=\frac{1}{\lambda}-\frac{1}{\lambda}e^{-\frac{\mathcal{L}_{\ell}}{\theta}\lambda},$ (5)
while the expected duration on a high fee pool is $d_{h}=\frac{1}{\lambda}.$ Thus,
@@
The expected duration of a liquidity cycle is shorter in the low fee pool than the high feel pool. Or, $d_{\ell}<d_{h}$.
Lemma 1 is intuitive. A liquidity cycle on the high fee pool only ends with the arrival of a large trader. Conversely, a liquidity cycle on the low fee pool ends either because a large liquidity taker arrived or the cumulative small liquidity trader orders exhaust the pool.
Notice that the low fee pool provides liquidity for a large share of the order flow, as both small and large LTs trade there. However, the liquidity providers on that pool earn a low fee per traded unit. Additionally, as the expected cycle duration is shorter, they need to manage their liquidity more often, which leads to larger gas costs per unit of time. Liquidity providers face a choice between the low and high fee pool or not participating in the market. Thus, an LPof size $q_{i}$ chooses between:
[…]
First, consider the choice between pools. Rearranging Equation 6, liquidity provider $i$ chooses the low-fee pool if and only if
$\left(d_{h}\ell-d_{\ell}h\right)q_{i}>\Gamma\left(d_{h}-d_{\ell}\right).$ (7)
Equation 7 highlights the tradeoff between the expected fee revenue per unit time and the fixed cost of accessing the market. If $\frac{h}{d_{h}}>\frac{\ell}{d_{\ell}}$, then the expected liquidity fee per unit of time on the high fee pool is larger than that of the low fee pool, and the left hand side of Equation 7 is negative. From Lemma 1, the expected duration is higher on the high fee pool and the right hand side is always positive. In this case, all liquidity providers choose the high fee pool. The more natural case to consider is if $\frac{h}{d_{h}}<\frac{\ell}{d_{\ell}}$ so that liquidity providers face a trade-off between a higher liquidity fee per unit of time on the low fee pool against lower gas costs on the high fee pool. Clearly, the larger a liquidity providers’ endowment, the more important is the fee revenue. We have
@@
For any liquidity pools, $\{\mathcal{L}_{\ell},\mathcal{L}_{h}\}$, for which $\frac{h}{d_{h}}<\frac{\ell}{d_{\ell}}$, if a liquidity provider of size $q$ prefers the low fee pool, then any liquidity provider with a larger endowment, $\widetilde{q}>q$, also prefers the low fee pool.
Now consider the choice of participating in the market. An agent only provides liquidity if she is able to break even on the high-fee pool – that is, if her endowment $q_{i}$ is large enough. The participation constraint follows from equation (6):
$q_{i}h-\Gamma\geq 0.$ (8)
Define $\underline{q}=\frac{\Gamma}{h}$. Recall, that the lower bound of the Pareto distribution is 1. Thus, if $\frac{\Gamma}{h}>1$, all LPs with $q_{i}>\underline{q}$ enter the market, and the marginal entrant earns zero expected profit. Conversely, if $\frac{\Gamma}{h}\leq 1$, liquidity provision is not competitive; all LPs enter the market and earn strictly positive profits.
Following Lemma 2, let $q_{\text{t}}$ be the threshold endowment such that all LPs with $q_{i}>q_{\text{t}}$ post liquidity on the low-fee pool and all LPs with $q_{i}\leq q_{\text{t}}$ choose the high-fee pool. This allows us to characterize pool sizes $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ as a function of the endowment for the marginal LP:
$\mathcal{L}_{\ell}$ $=\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i=\frac{Q}{Q-1}\left(\log Q-\log q_{\text{t}}\right)\text{ and }$
$\mathcal{L}_{h}$ $=\int_{\underline{q}}^{q_{\text{t}}}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i=\frac{Q}{Q-1}\left(\log q_{\text{t}}-\log\underline{q}\right)$ (9)
From equations (7) through (2.1) it follows that the expected profit difference between the low- and high-fee pools can be written as an increasing function of the marginal LP’s endowment, that is
@@
$\pi_{\ell}-\pi_{h}$ $=\frac{1}{\lambda d_{h}d_{\ell}}\left[\exp\left(-\frac{\lambda}{\theta}\mathcal{L}_{\ell}\right)\underbrace{\left(q_{i}h-\Gamma\right)}_{>0}-q_{i}\left(h-\ell\right)\right]$
@@
$=\frac{1}{\lambda d_{h}d_{\ell}}\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\underbrace{\left(q_{i}h-\Gamma\right)}_{>0}-q_{i}\left(h-\ell\right)\right].$ (10)
@@
    If $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}<1$ and $\frac{\Gamma}{h}<1-\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}$, then all LPs deposit liquidity on the low fee pool.
@@
    If $\Gamma>Q\ell$, then all LPs deposit liquidity on the high fee pool.
  3. iii.
[…]
Figure 2 illustrates the equilibrium regions in Proposition 1. If the gas price is low and LPs are homogeneous (low $\Gamma$ and $Q$), then all liquidity providers choose the low fee pool because managing liquidity is relatively cheap. If more high-endowment LPs enter the market (i.e., there is an increase in $Q$) the more liquidity is posted the low fee pool. Keeping the LT arrival rate fixed, a larger pool depletes more slowly and thus the liquidity cycle on the pool becomes longer. As a consequence, the liquidity fee per unit of time on pool $L$ drops and smaller LPs switch to the high-fee pool. As a result, liquidity becomes fragmented across the two pools.
@@
This figure plots the existence conditions for a fragmented market equilibrium, as described in Proposition 1, for various values of gas cost ($\Gamma$) on the y-axis and liquidity endowment $Q$ on the x-axis. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$ $\theta=0.66$.
An equilibrium in which all liquidity consolidates on the high fee pool is only sustainable for very high gas costs $\Gamma>Q\ell$. In this case, none of the LPs breaks even on pool $L$. For intermediate values of gas price, both pools co-exist with positive market share.
Proposition 2 establishes comparative statics for the two pools’ liquidity market shares. From equation (2.1), we can compute the liquidity market share of the low-fee pool at the beginning of each cycle as
$w_{\ell}=\frac{\mathcal{L}_{\ell}}{\mathcal{L}_{\ell}+\mathcal{L}_{h}}=\frac{\log Q-\log q_{\text{t}}^{\star}}{\log Q-\log\underline{q}}.$ (12)
@@
In equilibrium, the market share of the low fee pool $w_{\ell}$
  • i.
    decreases in the gas cost ($\Gamma$), the arrival rate of large trades ($\lambda$), and the fee on pool $H$ (h).
  • ii.
    increases in the fee on pool $L$ ($\ell$) and the arrival rate of small trades ($\theta$)
The results in Proposition 2 are intuitive. The market share of the low fee pool increases if the fee gap $h-\ell$ is narrower, since this reduces LP incentives to switch to the high-fee pool. If the small LT arrival rate is large, then liquidity cycles in the low-fee pool are shorter, increasing the revenue per unit of time and consequently the market share of pool $L$. Conversely, if large trades arrive more often (high $\lambda$), then the high fee pool attracts a higher share of incoming order flow and becomes more appealing for liquidity providers.
Figure 3 shows that the market share of the low fee pool (weakly) decreases in the gas cost $\Gamma$. A larger gas price increases the costs of active liquidity management, everything else equal, and incentivizes smaller LPs to switch from the low fee pool to the high fee pool, since the latter has a lower turnover. For $\Gamma\leq h$, any increase in gas costs leads to a redistribution of liquidity from one pool to another; the aggregate liquidity across both pools is constant since all LPs participate in the market.
@@
This figure illustrates the equilibrium liquidity market shares (left panel) and absolute liquidity deposits on the two pools (right panel), as a function of the gas fee $\Gamma$. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
If gas prices increase beyond a threshold ($\Gamma>h$), then the aggregate liquidity falls since LPs with $q_{i}<\frac{\Gamma}{h}$ are shut out of the market. Both the low and high fee pool experience a decrease in liquidity deposits. However, the liquidity drop is sharper for the low fee pool which further depresses its market share.

### v7

@@ 2.1 Equilibrium
2.1.1 Optimal trade size
First, consider the decisions of arbitrageurs and liquidity traders holding a value $v\left(1+\delta\right)$ for the asset. Faced with pool sizes of $T_{\ell}$ and $T_{h}$ in the low and high pool respectively, their optimal trade on pool $k$ maximizes their expected profit, net of fees and price impact:
$\max_{\tau}\text{Profit $\textbf{LT}$}\left(\tau,\delta\right)\equiv\tau v\left(1+\delta\right)-\left(1+f_{k}\right)\tau T_{k}\frac{v\left(1+r\right)}{\tau+\left(1+r\right)\left(T_{k}-\tau\right)},$ (6)
which yields the optimal trade quantity:
$\tau^{\star}\left(\delta\right)=T_{k}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f_{k}}{1+\delta}}\right\}\right\}.$ (7)
From equation (7), a trader with valuation $v\left(1+\delta\right)$ only trades on pool $k$ if the gains from trade are larger than the liquidity fee, i.e., $\delta>f_{k}$. Further, if $\delta>\left(1+f_{k}\right)\left(1+r\right)^{2}-1$ so that the gains from trade are larger than the maximum price impact, then the trader consumes all available liquidity in the pool.
2.1.2 Fee revenue for liquidity providers from private value trades
The revenue for liquidity providers can be expressed as the product of the pool fee and the numéraire deposit required from liquidity traders to purchase $\tau^{\star}$ token units, denoted by $n(\tau^{\star},T_{k})$. That is,
$f_{k}n\left(\tau^{\star}\left(\delta\right),T_{k}\right)=f_{k}vT_{k}\min\left\{1+r,\frac{1+r}{r}\max\left\{0,\sqrt{\frac{1+\delta}{1+f_{k}}}-1\right\}\right\}.$ (8)
If the innovation $\delta$ corresponds to a private rather than common value shock, then an arbitrageur optimally steps in to reverse the liquidity trade as described in Lehar and Parlour (forthcoming). In this case, liquidity providers effectively earn double the fee revenue in (8) without affecting the capital structure of the pool; there is neither a capital gain nor a loss for the LPs.
The fee revenue in (8) scales linearly with the size of the pool $T_{k}$. Since fee proceeds are distributed pro-rata among liquidity providers based on their share $\frac{q_{i}}{T_{k}}$, it follows that fee revenue an LP with endowment $q_{i}$ providing liquidity on pool $k$ increases linearly in their endowment:
$\text{FeeRevenue}_{i,k}\left(\delta\right)$ $=2\frac{q_{i}}{T_{k}}f_{k}n\left(\tau^{\star},T_{k}\right)$
$=2q_{i}vf_{k}\min\left\{1+r,\frac{1+r}{r}\max\left\{0,\sqrt{\frac{1+\delta}{1+f_{k}}}-1\right\}\right\}.$ (9)
The expression in (2.1.2) denotes the fee revenue conditional on the private value $\delta$ of the incoming trade. To compute the expected fee revenue, we integrate this expression across all posible value shocks:
$\mathbb{E}\text{FeeRevenue}_{i,k}$ $=\int_{\delta=1}^{\Delta^{2}-1}\text{FeeRevenue}_{i,k}\left(\delta\right)\phi\left(\delta\right)\mathop{}\!\mathrm{d}\delta$
$=q_{i}\underbrace{v\frac{f_{k}(r+1)\left(2\Delta-r\sqrt{f_{k}+1}-2\sqrt{f_{k}+1}\right)}{\Delta}}_{\equiv\mathcal{L}\left(f_{k}\right)},$ (10)
where we define $\mathcal{L}\left(f_{k}\right)$ as the liquidity yield: that is, the per-unit LP fee revenue from supplying liquidity to LTs in pool $k$.
@@ 2.1.2 Fee revenue for liquidity providers from private value trades
There exists a threshold fee level $\overline{f}>0$ such that the liquidity revenue $\mathcal{L}\left(f_{k}\right)$ first increases in the pool fee $f_{k}$ for $f\leq\overline{f}$, then decreases in the pool fee for $f>\overline{f}$.
Lemma 1 points out to a non-linear relationship between fee levels and liquidity yield. Initially, as fees increase, the enhanced revenue from higher fees outweighs the decrease in trading volume due to increased transaction costs, resulting in a net gain in revenue. However, beyond a certain fee threshold, the drop in trading volume dominates the larger fee, leading to a decrease in overall revenue. A salient implication is that if pool fees are large enough, the liquidity yield on the high fee pool may exceed the yield on the low-fee pool.
2.1.3 Adverse selection cost for liquidity providers
If news occurs (i.e., if $\delta$ represents a common value shock), liquidity providers trade against arbitrageurs rather than liquidity traders. In this case, there is no subsequent price reversal following the initial trade. The capital structure of the liquidity pool changes, as arbitrageurs remove the more valuable asset: i.e., buy tokens upon a positive common value shock. While LPs earn fee revenues on arbitrage trades, they also incur adverse selection losses by trading against the direction of the news. Moreover, if the magnitude of news is large enough that arbitrageurs remove all tokens supplied in the price range, then LPs face additional costs, that is a gas fee $\Gamma\mathop{}\!\mathrm{d}i$ to re-balance liquidity around the new asset value.
Table 1 delineates the LP fee revenue from selling tokens to arbitrageurs, as well as the marked-to-market value of the tokens sold. If the size of news ($\delta$) does not exceed the pool fee, then arbitrageurs do not trade since the potential profit does not justify the transaction cost. Conversely, if the news size is larger than pool fee, then arbitrageurs execute a trade proportional to the size of the pool, and they exhaust the available liquidity on the price range if the news is large enough: specifically, if $\delta>\left(1+f_{k}\right)\left(1+r\right)^{2}-1$. The profit for liquidity providers in each scenario is the difference between the revenue and the marked-to-market value. Notably, the profit is consistently negative, since LPs are trading against the direction of news.
Table 1: Fee revenue and capital losses on arbitrage trades
News size Revenue (numeraire) Marked-to-market token value
$\delta\leq f_{k}$ 0 0
$\delta\in\left(f_{k},\left(1+f_{k}\right)\left(1+r\right)^{2}-1\right]$ $vq_{i}\frac{1+r}{r}\left(1+f_{k}\right)\left(\sqrt{\frac{1+\delta}{1+f_{k}}}-1\right)$ $vq_{i}\frac{1+r}{r}\left(1+\delta\right)\left(1-\sqrt{\frac{1+f_{k}}{1+\delta}}\right)$
$\delta>\left(1+f_{k}\right)\left(1+r\right)^{2}-1$ $vq_{i}\left(1+f_{k}\right)\left(1+r\right)$ $vq_{i}\left(1+\delta\right)$
The expected LP profit from trading with arbitrageurs equals $-q_{i}v\times\mathcal{A}\left(f_{k}\right)$, where $\mathcal{A}\left(f_{k}\right)$ is the per-unit adverse selection cost from liquidity provision in pool $k$:
$\mathcal{A}\left(f_{k}\right)$ $=\mathbb{P}\left(f_{k}<\delta\leq(1+f_{k})(1+r)^{2}-1\right)\times\frac{1+r}{r}\mathbb{E}\left[\left(1+f_{k}\right)+\left(1+\delta\right)-2\sqrt{\left(1+\delta\right)\left(1+f_{k}\right)}\right]+$
$+\mathbb{P}\left(\delta>(1+f_{k})(1+r)^{2}-1\right)\times\left[\mathbb{E}\left(1+\delta\right)\right]-\left(1+f_{k}\right)\left(1+r\right)\Big{\}}.$ (11)
@@ 2.1.3 Adverse selection cost for liquidity providers
The adverse selection cost $\mathcal{A}\left(f_{k}\right)$ decreases in the pool fee $f_{k}$. In particular, the high-fee pool has a lower adverse selection cost than the low-fee pool.
Lemma 2 indicates that higher pool fees lower adverse selection costs through two mechanisms: First, they increase compensation per unit traded for liquidity providers (LPs), enhancing their returns on trades with arbitrageurs. Second, higher fees discourage arbitrageur activity, effectively reducing the volume of informed trades. Figure 2 showcases the results in Lemmas 1 and 2 and illustrates the comparative statics of liquidity yield and adverse selection cost with respect to the pool fee.
Figure 2: Liquidity yield and adverse selection cost
This figure illustrates the expected fee yield from liquidity trades (left panel) and the adverse selection cost (right panel), as a function of the pool fee $f$. Parameter values: $r=0.001$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
Liquidity rebalancing costs arise only when news events are large enough to deplete all available liquidity within a given price range, pushing liquidity providers’ (LPs) positions “out of range.” Rebalancing only occurs post-news, since equally large liquidity trades would be reversed by arbitrageurs. Conditional on news arrival, the expected cost of rebalancing is
@@
$\mathcal{C}\left(k\right)=\mathbb{P}\left(\delta>(1+f_{k})(1+r)^{2}-1\right)\Gamma=\Gamma\left(1-\frac{\sqrt{1+f_{k}}\left(1+r\right)}{\Delta}\right),$ (12)
which is decreasing in the pool fee $f_{k}$. This result is straightforward: smaller news events can cause arbitrageurs to deplete liquidity in low-fee pools, whereas it takes larger news to do the same in high-fee pools. Consequently, LPs in lower fee pools face more frequent rebalancing and incur higher fixed costs.
2.1.4 Liquidity provider pool choice
Liquidity providers face a choice between the low and high fee pool or not participating in the market. An LP of size $q_{i}$ earns expected profit
@@ 2.1.4 Liquidity provider pool choice
$\pi_{L}$ $=q_{i}\left[\left(1-\eta\right)\mathcal{L}\left(\ell\right)-\eta\mathcal{A}\left(\ell\right)\right]-\eta\Gamma\left(1-\frac{\sqrt{1+\ell}\left(1+r\right)}{\Delta}\right)\text{ and }$ (13)
$\pi_{H}$ $=q_{i}\left[\left(1-\eta\right)\mathcal{L}\left(h\right)-\eta\mathcal{A}\left(h\right)\right]-\eta\Gamma\left(1-\frac{\sqrt{1+h}\left(1+r\right)}{\Delta}\right),$
from choosing pool $L$ or $H$, respectively. Equation (13) underscores the trade-off faced by liquidity providers (LPs): balancing the liquidity yield from trades with liquidity traders (LTs) against the adverse selection costs and the fixed gas costs associated with re-balancing their position.
First, consider the choice of participating in the market. An agent only provides liquidity on pool $k$ if she is able to break even – that is, if her endowment $q_{i}$ is large enough. We define $\underline{q}_{L}$ and $\underline{q}_{H}$ as the thresholds at which the participation constraints $\pi_{L}\left(q\right)=0$ and $\pi_{H}\left(q\right)=0$ are satisfied, respectively. If $\underline{q}_{k}\geq 0$ for a pool $k$, it indicates that any LP with an endowment $q_{i}$ at least equal to $\underline{q}_{k}$ can join pool $k$ and expect to earn a positive profit, with the marginal entrant breaking even. Conversely, if $\underline{q}_{k}<0$, it suggests that pool $k$ is not economically viable as the participation constraint is breached for all LPs.
Assumption 2:
To avoid trivial cases, we focus on the case that both markets are potentially viable, or equivalently the intensity of news is low enough:
$\eta\leq\min_{k}\frac{\mathcal{L}\left(k\right)}{\mathcal{L}\left(k\right)+\mathcal{A}\left(k\right)},$ (14)
such that $\underline{q}_{k}\geq 0$.
Next, consider the choice between pools. Liquidity provider $i$ chooses the low-fee pool if and only if
$\pi_{L}-\pi_{H}=q_{i}\left[\left(1-\eta\right)\left(\mathcal{L}\left(\ell\right)-\mathcal{L}\left(h\right)\right)+\eta\underbrace{\left(\mathcal{A}\left(h\right)-\mathcal{A}\left(\ell\right)\right)}_{<0}\right]-\underbrace{\Gamma\frac{\eta(1+r)}{\Delta}\left(\sqrt{1+h}-\sqrt{1+\ell}\right)}_{>0}>0.$ (15)
Liquidity providers in the high-fee pool face both lower adverse selection and rebalancing costs. Therefore, the low-fee pool can only be chosen in equilibrium if it offers a higher liquidity yield, specifically if $\mathcal{L}(\ell)-\mathcal{L}(h)>0$, and if the intensity of news $\eta$ is sufficiently low. Otherwise, all liquidity providers prefer to supply tokens to the high-fee pool if the participation constraint is satisfied. Further, equation (15) highlights the economies of scale embedded in liquidity provision with fixed rebalancing costs. That is, if a liquidity provider of size $q$ prefers the low fee pool, then any liquidity provider with a larger endowment, $\widetilde{q}>q$, also prefers the low fee pool.
@@
    If $\eta>\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$, then all LPs with $q_{i}>\underline{q}_{h}$ deposit liquidity on the high fee pool.
@@
    Otherwise, there exists a unique fragmented equilibrium characterized by marginal trader $q_{\text{t}}^{\star}>\underline{q}_{h}$ which solves
    $q_{\text{t}}^{\star}=\Gamma\frac{\eta(1+r\left(\sqrt{1+h}-\sqrt{1+\ell}\right))}{\Delta\left[\left(1-\eta\right)\left(\mathcal{L}\left(\ell\right)-\mathcal{L}\left(h\right)\right)+\eta\left(\mathcal{A}\left(h\right)-\mathcal{A}\left(\ell\right)\right)\right]}$ (16)
    such that all LPs with $q_{i}\in\left(\underline{q}_{h},q_{\text{t}}^{\star}\right]$ deposit liquidity in the high fee pool and all LPs with $q_{i}>q_{\text{t}}^{\star}$ choose the low fee pool.
Figure 3 illustrates the equilibrium regions in Proposition 1. When news intensity $\eta$ is high, or pool $H$ offers a substantially higher fee than pool $L$, liquidity suppliers gravitate towards pool $H$, resulting in a single-maker equilibrium. Conversely, a lower $\eta$ translates to lower adverse selection costs. If this is the case, or if the fee differential between the two pools is low, liquidity providers with large endowments $q$ migrate to the lower-fee pool to compete for order flow from small traders, causing liquidity to fragment between the two pools.
@@
This figure plots the existence conditions for a fragmented market equilibrium, as described in Proposition 1, for various values of the news intensity ($\eta$) on the y-axis and liquidity fee on pool $L$ on the x-axis. Parameter values: $r=0.001$, $h=2$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
Proposition 2 establishes the impact of gas prices on the two pools’ liquidity market shares. We can compute the liquidity market share of the low-fee pool in a fragmented equilibrium as
$w_{\ell}=\frac{\exp\left(-\frac{q_{\text{t}}-\underline{q}_{h}}{\lambda}\right)\left(q_{\text{t}}+\lambda\right)}{\underline{q}_{h}+\lambda}\leq 1,$ (17)
with equality for $\Gamma=0$. That is, as fixed costs drop to zero, the low fee pool asymptotically captures the full market share.
@@
In equilibrium, the market share of the low fee pool $w_{\ell}$ decreases in the gas cost ($\Gamma$).
We stress the critical role of fixed gas costs in driving market fragmentation. Since liquidity fee revenues and adverse selection costs are distributed pro-rata, in the absence of gas fees, all liquidity providers (LPs) would converge on a single pool — the one offering the optimal balance between fee yield and informational costs. For instance, if $\Gamma=0$, all LPs would select pool $L$ if the news arrival rate is sufficiently low, as defined by $\eta\leq\frac{\mathcal{L}(l)-\mathcal{L}(h)}{\mathcal{L}(l)-\mathcal{L}(h)+\mathcal{A}(l)-\mathcal{A}(h)}$, or choose pool $H$ otherwise. It is the introduction of fixed costs that drives LPs to segregate into different pools based on their size.
Figure 4 shows that the market share of the low fee pool decreases in the gas cost $\Gamma$. A larger gas price increases the costs of re-balancing upon the arrival of large enough news, everything else equal, and incentivizes smaller LPs to switch from the low fee pool to the high fee pool, since the arbitrageurs are less likely to fully consume liquidity there. Further, the right panel illustrates the extensive margin effect of gas prices: any increase in gas costs leads to a decrease in aggregate liquidity supply as some LP with low endowments are driven out of the market (that is, the threshold $\underline{q}_{h}$ increases in $\Gamma$).
@@
This figure illustrates the equilibrium liquidity market shares (left panel) and the aggregate liquidity supply on the two pools (right panel), as a function of the gas fee $\Gamma$. Parameter values: $r=0.001$, $h=2$, $\ell=1$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.